# Efficient4D: Fast Dynamic 3D Object Generation from a Single-view Video

### v4

@@ 5.1 Experiment setup
[…]
Competitors For comparison, we mainly focus on video-to-4D and image-to-4D task. The competitors include Consistent4D [14] (ICLR2024), 4DGen [18] (ArXiv2024) and STAG4D [19] (ECCV2024) for video-to-4D, and Animate124 [52] (ArXiv2023) and DreamGaussian4D [53] (ArXiv2024) for image-to-4D. We obtained their results by running their released official code. We also compare SyncDreamer [8] partially by replacing SyncDreamer-T in ablation study (Section 5.6).
[…]
Evaluation data To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. For video-to-4D, we focused on 36 sequences: 32 sequences released by [14] and 4 sequences processed by ourselves. Among the released data, seven sequences are synthetic data where ground truth are available. For image-to-4D, we used the ten images released by [52]. Our four sequences are used for sparse input evaluation in Section 5.5, which only contain two frames for each sequence. Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [54]. The other one named yellow face was collected from the internet. All the data are publicly available.
[…]
Evaluation metrics As 4D generation research is still at early stage, there is no well established metric yet. However, we adopt multiple metrics by referring related works for comprehensive evaluation. For evaluating on synthetic data, we use LPIPS score [55] and CLIP similarity [56] between rendered images and ground truth following [14]. For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [57, 53] to measure image quality. For temporal smoothness, following [18, 58, 59] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views. To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
[…]
Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase. In practice, we can adjust the weight following two rules: (i) the middle weight should be at least 0.50.50.5; (ii) At least half of the volumes should have a positive weight. Here we provide some sample weights which have a good effect:
@@
where $N$ denotes the frame number. The weight will be normalized to ensure the sum is 111.
In denosing process, we follow the default setting of [8] using a improved sampling strategy introduced by HarmonyView [60]. HarmonyView redefines the score estimation by decomposing consistency and diversity. Please refer to [60] for details.
In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 50,000 without densification. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 500 iterations with batch size 1. All other 4Dgs hyperparameters remain consistent with those in [15]. The balancing weights in loss function (equation 13) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$. In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [50]. For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
@@ 5.2 Evaluation on synthetic data
Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
@@
Consistent4D [14] 4DGen [18] STAG4D [19] Efficient4D(Ours)
@@
We first present the results on synthetic data in Figure 4 with ground truth shown. Since the data only include videos, we compare with two video-to-4D methods: Consistent4D [14], 4DGen [18] and STAG4D [19]. Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth, such as the direction of dinosaur’s head, the color of trump’s arm and the teeth of skull. The quantitative metrics are also presented in Table 1. The better CLIP and LPIPS scores further validate our superior generation results.
@@ 5.3 Video-to-4D comparison
Figure 5: Qualitative comparisons on video-to-4D generation. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19]. For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
Figure 6: Qualitative comparisons on image-to-4D generation. We compare our Efficient4D with Animate124 [52] and DreamGaussian4D [53] (DG4D). For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
@@
Consistent4D [14] 0.8471 0.9692 0.9658 0.9697 120 mins
4DGen [18] 0.8730 0.9568 0.9568 0.9573 130 mins
STAG4D [19] 0.8398 0.9766 0.9731 0.9760 70 mins
@@
Animate124 [52] 0.8076 0.9673 0.9639 0.9541 540 mins
DreamGaussian4D [53] 0.8242 0.8999 0.9072 0.9038 12 mins
@@
In Table 2, we compare our method with baselines quantitatively. The metrics and data used are described in Section 5.1. We draw several observations: (i) Our method achieves superior image quality and temporal smoothness on all cases validated by the higher CLIP and CLIP-T scores; (ii) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement (10 mins vs 120 mins). More specifically, the speed improvement is attributed to (i) our image supervision design, converging optimization faster and reduce the training iterations, and (ii) each iteration in our method requiring much less time thanks to the efficiency of Gaussian representation.
@@ 5.5 Challenging cases
Input Ours Consistent4D [14]
@@
We assessed our method’s performance on two challenging cases: extremely sparse input and rotation dynamics. In sparse case, the input comprising only two discrete frames. In such cases, we set $w=0.25$ in equation (9). We also modified the code of Consistent4D [14] to make it adapt to such case. As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions. Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
For another rotation case, we adopt the data introduced in [61]. As shown in the first row of Figure 8, the character will rotate its body in the monocular video input, which is complex and challenging. However, our approach can also successfully infer the accurate geometry while Consistent4D tends to generate fragments. For example, from the Consistent4D results in Figure 8, we observe that the punching figure and the kicking man both suffer fragmented arms or legs.
@@ 5.6.1 Ablation of image generation
Figure 10: Ablation study on the temporal synchronization. We compare SV3D [62], SyncDreamer [8], SyncDreamer with consistent self-attention [63] (SyncDreamer-A) and our proposed Syncdreamer-T. We show nine consecutive frames for each case to better evalute the temporal consistency.
@@
[…]
Ablation on temporal synchronization We also examine the design of temporal synchronization. In conceptual terms, directly applying image-to-3D methods to each frame independently is inferior due to intrinsic generative randomness. This can be validated in Figure 10 by comparing SV3D [62], SyncDreamer [8] and SyncDremaer-T(Ours). The image-to-3D methods (SV3D and SyncDreamer) cannot generate a consistent shape or texture across frames. Therefore, a proper temporal synchronization mechanism is necessary. In Section 4.1, we propose SyncDreamer-T, which blends the 3D volume feature. We also implement a consistent self-attention mechanism introduced in [63] in SyncDreamer [8], resulting in Syncdreamer-A. Specifically, in each self-attention block that projects $Q_{i},K_{i},V_{i}$ corresponding to the image $I_{i}$, we randomly sample another image $I_{j}$’s $K_{j},V_{j}$, and concatenate two $K$s and $V$s as the new $K_{i}$ and $V_{i}$. The results in Figure 10 indicate that our method generates a more consistent texture and smoother motion than SyncDreamer-A. We ascribe this to the following reason: The cross-attended features may focus too much on the details that its control on global texture and motion is weakened.
@@ 5.6.2 Ablation of reconstruction stage
Figure 13: Ablation study on the choice of 4D representation. We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF [14], Deformable-3DGS [44] and 4DGS [15]. 4DGS features the fastest convergence speed.
In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, (3) SDS loss, and (4) choice of 4D representation. The visual results are shown in Figure 11, 12 and 13.
[…]
Effect of confidence map The integration of a confidence-weighted loss in our design serves as a strategy to mitigate training data noise. In the left of Figure 11, we study the effect of confidence map when obvious inconsistency exists in the generated images. Our temporal smooth mechanism can support consistency for most areas, whilst a small fraction of lower-quality regions may introduce conflicting gradients in the reconstruction stage, thus hurting the overall quality. To mitigate this, our confidence aware design comes into play for weakening the supervision from those inconsistent regions. Since these weakly supervised regions only occupy a small proportion across all views, the missing information about texture or geometry can be compensated by the redundancy of other views and generalization ability of 4D Gaussians. As shown in the left of Figure 11, the inclusion of confidence maps effectively reduces blurry and inconsistent rendering even when the generated images have plausible consistency, resulting in a significant overall improvement in quality and temporal smoothness.
[…]
@@
[…]
Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF [14] and Deformable-3DGS [44] used in other 4D generation works [14, 53]. In the comparison shown in Figure 13, 4DGS [15] converge in 500 iterations while the others do not. Also, thanks to the fast convergence speed of 4DGS, our method achieves significant efficiency advantage.

### v5

@@ 5.1 Experiment setup
Competitors For comparison, we mainly focus on video-to-4D and image-to-4D task. The competitors include Consistent4D Jiang et al. (2024) (ICLR2024), 4DGen Yin et al. (2023) (ArXiv2024) and STAG4D Zeng et al. (2024) (ECCV2024) for video-to-4D, and Animate124 Zhao et al. (2023) (ArXiv2023) and DreamGaussian4D Ren et al. (2023) (ArXiv2024) for image-to-4D. We obtained their results by running their released official code. We also compare SyncDreamer Liu et al. (2024) partially by replacing SyncDreamer-T in ablation study (Section 5.6).
Evaluation data To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. For video-to-4D, we focused on 36 sequences: 32 sequences released by Jiang et al. (2024) and 4 sequences processed by ourselves. Among the released data, seven sequences are synthetic data where ground truth are available. For image-to-4D, we used the ten images released by Zhao et al. (2023). Our four sequences are used for sparse input evaluation in Section 5.5, which only contain two frames for each sequence. Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab 40. The other one named yellow face was collected from the internet. All the data are publicly available.
Evaluation metrics As 4D generation research is still at early stage, there is no well established metric yet. However, we adopt multiple metrics by referring related works for comprehensive evaluation. For evaluating on synthetic data, we use LPIPS score Zhang et al. (2018) and CLIP similarity Radford et al. (2021) between rendered images and ground truth following Jiang et al. (2024). For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as Tang et al. (2024); Ren et al. (2023) to measure image quality. For temporal smoothness, following Yin et al. (2023); Esser et al. (2023); Geyer et al. (2023) we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views. To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
Implementation details In equation (9), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase. In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight. Here we provide some sample weights which have a good effect:
@@
where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
In denosing process, we follow the default setting of Liu et al. (2024) using a improved sampling strategy introduced by HarmonyView Woo et al. (2024). HarmonyView redefines the score estimation by decomposing consistency and diversity. Please refer to Woo et al. (2024) for details.
In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 50,000 without densification. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 500 iterations with batch size 1. All other 4DGS hyperparameters remain consistent with those in Yang et al. (2024a). The balancing weights in loss function (equation 14) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$. In equation (13), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model Huang et al. (2022). For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
@@ 5.2 Evaluation on synthetic data
Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D Jiang et al. (2024), 4DGen Yin et al. (2023) and STAG4D Zeng et al. (2024).
@@
Consistent4D 4DGen STAG4D Efficient4D(Ours)
@@
We first present the results on synthetic data in Figure 4 with ground truth shown. Since the data only include videos, we compare with two video-to-4D methods: Consistent4D Jiang et al. (2024), 4DGen Yin et al. (2023) and STAG4D Zeng et al. (2024). Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth, such as the direction of dinosaur’s head, the color of trump’s arm and the teeth of skull. The quantitative metrics are also presented in Table 1. The better CLIP and LPIPS scores further validate our superior generation results.
@@ 5.3 Video-to-4D comparison
Figure 5: Qualitative comparisons on video-to-4D generation. We compare our Efficient4D with Consistent4D Jiang et al. (2024), 4DGen Yin et al. (2023) and STAG4D Zeng et al. (2024). For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
Figure 6: Qualitative comparisons on image-to-4D generation. We compare our Efficient4D with Animate124 Zhao et al. (2023) and DreamGaussian4D Ren et al. (2023) (DG4D). For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
@@
Consistent4D Jiang et al. (2024) 0.8471 0.9692 0.9658 0.9697 120 mins
4DGen Yin et al. (2023) 0.8730 0.9568 0.9568 0.9573 130 mins
STAG4D Zeng et al. (2024) 0.8398 0.9766 0.9731 0.9760 70 mins
@@
Animate124 Zhao et al. (2023) 0.8076 0.9673 0.9639 0.9541 540 mins
DreamGaussian4D Ren et al. (2023) 0.8242 0.8999 0.9072 0.9038 12 mins
@@
[…]
@@ 5.5 Challenging cases
Input Ours Consistent4D
@@
We assessed our method’s performance on two challenging cases: extremely sparse input and rotation dynamics. In sparse case, the input comprising only two discrete frames. In such cases, we set $w=0.25$ in equation (10). We also modified the code of Consistent4D Jiang et al. (2024) to make it adapt to such case. As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions. Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
For another rotation case, we adopt the data introduced in Yang et al. (2025). As shown in the first row of Figure 8, the character will rotate its body in the monocular video input, which is complex and challenging. However, our approach can also successfully infer the accurate geometry while Consistent4D tends to generate fragments. For example, from the Consistent4D results in Figure 8, we observe that the punching figure and the kicking man both suffer fragmented arms or legs.
@@ 5.6.1 Ablation of image generation
Ablation on temporal synchronization We also examine the design of temporal synchronization. In conceptual terms, directly applying image-to-3D methods to each frame independently is inferior due to intrinsic generative randomness. This can be validated in Figure 13 by comparing SV3D Voleti et al. (2024), SyncDreamer Liu et al. (2024) and SyncDremaer-T(Ours). The image-to-3D methods (SV3D and SyncDreamer) cannot generate a consistent shape or texture across frames. Therefore, a proper temporal synchronization mechanism is necessary. In Section 4.1, we propose Sync-Dreamer-T, which blends the 3D volume feature. We also implement a consistent self-attention mechanism introduced in Zhou et al. (2024) in SyncDreamer Liu et al. (2024), resulting in Syncdreamer-A. Specifically, in each self-attention block that projects $Q_{i},K_{i},V_{i}$ corresponding to the image $I_{i}$, we randomly sample another image $I_{j}$’s $K_{j},V_{j}$, and concatenate two $K$s and $V$s as the new $K_{i}$ and $V_{i}$. The results in Figure 13 indicate that our method generates a more consistent texture and smoother motion than SyncDreamer-A. We ascribe this to the following reason: The cross-attended features may focus too much on the details that its control on global texture and motion is weakened.
@@ 5.6.2 Ablation of reconstruction stage
Figure 12: Ablation study on the choice of 4D representation. We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF Jiang et al. (2024), Deformable-3DGS Wu et al. (2024a) and 4DGS Yang et al. (2024a). 4DGS features the fastest convergence speed.
In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, (3) SDS loss, and (4) choice of 4D representation. The visual results are shown in Figure 10, 11 and 12.
Effect of confidence map The integration of a confidence weighted loss in our design serves as a strategy to mitigate training data noise. In the left of Figure 10, we study the effect of confidence map when obvious inconsistency exists in the generated images. Our temporal smooth mechanism can support consistency for most areas, whilst a small fraction of lower-quality regions may introduce conflicting gradients in the reconstruction stage, thus hurting the overall quality. To mitigate this, our confidence aware design comes into play for weakening the supervision from those inconsistent regions. Since these weakly supervised regions only occupy a small proportion across all views, the missing information about texture or geometry can be compensated by the redundancy of other views and generalization ability of 4D Gaussians. As shown in the left of Figure 10, the inclusion of confidence maps effectively reduces blurry and inconsistent rendering even when the generated images have plausible consistency, resulting in a significant overall improvement in quality and temporal smoothness.
@@
Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF Jiang et al. (2024) and Deformable-3DGS Wu et al. (2024a) used in other 4D generation works Jiang et al. (2024); Ren et al. (2023). In the comparison shown in Figure 12, 4DGS Yang et al. (2024a) converge in 500 iterations while the others do not. Also, thanks to the fast convergence speed of 4DGS, our method achieves significant efficiency advantage.
Figure 13: Ablation study on the temporal synchronization. We compare SV3D Voleti et al. (2024), SyncDreamer Liu et al. (2024), SyncDreamer with consistent self-attention Zhou et al. (2024) (SyncDreamer-A) and our proposed Syncdreamer-T. We show nine consecutive frames for each case to better evalute the temporal consistency.
@@ 6 Conclusion
7 Data Availability Statement
The datasets used for evaluation and comparison are available at Consistent4D Jiang et al. (2024) (https://github.com/yanqinJiang/Consistent4D), Animate124 Zhao et al. (2023) (https://github.com/HeliosZhao/Animate124/tree/main/benchmark) and Sketchfab 40 (https://sketchfab.com/).